OpenStudy (anonymous):

Mr. Quick bought a new computer system. The regular price was $2380, but he got a 15% discount. How much did he pay?

OpenStudy (anonymous):

357

OpenStudy (anonymous):

2380*(1.00-.15) = answer roundabouts 2000

OpenStudy (anonymous):

357 you multiply 2380 and .15

OpenStudy (anonymous):

just for a logical argument, if i got a 1% discount on a 1000 t.v., i would only have to pay $10 for that t.v.?? I REALLY HOPE MY NEAREST T.V. DEALER HAS SOME SWEET DEALS SOON!!! :D

OpenStudy (anonymous):

i dont get it, because on the bottom of my homework.. it has like a puzzle thing with the answers.. and the answers left are.. $1094, $2023, $454

OpenStudy (anonymous):

well you should tell your teacher and explain your logic and see what she/he says

OpenStudy (anonymous):

it's $2023, i just said "roundabouts 2000" cause i didn't know the exact answer. Think about it this way: The regular price is $2380. If you were to pay full price, you would pay 100% of the price. However, if you were given a 15% discount (which is awesome), now you get to pay less than 100%. 15% less, to be exact. 100-15 = 85% is what you are paying. So, you are paying 2380 * .85 (.85 is 85% in decimal form) 2380 *.85 = 2023!

OpenStudy (anonymous):

ohhh... i see, sorrry... lol

OpenStudy (anonymous):

$2380 * 85% = $2023